if y varies directly as x, and y=9 when x=4, find y when x=16
Answer by nerdybill(7384)

if y varies directly as x, and y=9 when x=4, find y when x=16
.
When you see "direct variation" think:
y = kx
.
To find k, use "y=9 when x=4":
y = kx
9 = k(4)
9/4 = k
.
so, our equation is:
y = (9/4)x
.
now, we can answer:
find y when x=16
y = (9/4)x
y = (9/4)(16)
y = (9)(4)
y = 36